OpenStudy (anonymous):

The rms speed of hydrogen molecules (h2) at 27°C (300 K) was calculated to be 1930 m/s. At what temperature would hydrogen molecules have an rms speed of 1.1x10^4 m/s, which is sufficient to escape from the earth? ?? little help please what equation should i use?

OpenStudy (anonymous):

Hi, you just need to know that the mean square speed of the molecules is proportional to the absolute temperature. The first piece of data lets you calculate the constant of proportionality. Note that mean square is not the same as root mean square.

OpenStudy (anonymous):

v=((3RT)/M)^1/2

OpenStudy (anonymous):

No i dont know what the mean square is just the rms speed, do i use this equation to find the constant of proportionality ? \[Vrms = \sqrt{(3KBT)/m}\]

OpenStudy (anonymous):

replace R=(KB)N then get KB replace yr equation

OpenStudy (anonymous):

mean square is not the rms

OpenStudy (anonymous):

Vrms=(V^2)^1/2

OpenStudy (anonymous):

u can use my equation above to solve it

OpenStudy (anonymous):

if you know the root mean square speed, then just square it to get the mean square speed !

OpenStudy (anonymous):

Then use the proportionality of temperature and mean square speed\[T=constant <v^2> \] The values of v squared and T at 300 K let you find the constant, then use that information to find T when the mean square speed is (1.1 * 10^4) squared.

OpenStudy (anonymous):

What sign is used for the constant in a equation? R?

OpenStudy (anonymous):

The constant that i have to calculate with this equation: constant = T/v^2 and then have to use the outcome in v=((3RT)/M)^1/2

OpenStudy (anonymous):

don't get it sorry haha..